### v1

# A Perspective on AI-Guided Molecular Simulations in VR: Exploring Strategies for Imitation Learning in Hyperdimensional Molecular Systems

Abstract
Molecular dynamics (MD) simulations are a crucial computational tool for researchers to understand and engineer molecular structure and function in areas such as drug discovery, protein engineering, and material design. Despite their utility, MD simulations are expensive, owing to the high dimensionality of molecular systems. Interactive molecular dynamics in virtual reality (iMD-VR) has recently been developed as a ‘human-in-the-loop’ strategy, which leverages high-performance computing to accelerate the researcher’s ability to solve the hyperdimensional sampling problem. By providing an immersive 3D environment that enables visualization and manipulation of real-time molecular motion, iMD-VR enables researchers and students to efficiently and intuitively explore and navigate these complex, high-dimensional systems. iMD-VR platforms offer a unique opportunity to quickly generate rich datasets that capture human experts’ spatial insight regarding molecular structure and function. This paper explores the possibility of employing user-generated iMD-VR datasets to train AI agents via imitation learning (IL). IL is an important technique in robotics that enables agents to mimic complex behaviors from expert demonstrations, thus circumventing the need for explicit programming or intricate reward design. We review the utilization of IL for manipulation task domains in robotics and discuss how iMD-VR recordings could be used as datasets to train IL models for interacting with MD simulations and solving specific molecular ‘tasks’. We then investigate how such approaches could be applied to the data structures captured from iMD-VR recordings. Finally, we outline the future research directions and potential challenges of using AI agents to augment human expertise to efficiently navigate vast conformational spaces, highlighting how this approach could provide valuable insight across domains such as materials science, protein engineering, and computer-aided drug design.
(Accepted for presentation at the First Workshop on "eXtended Reality & Intelligent Agents" (XRIA24) @ ECAI24, Santiago De Compostela (Spain), 20 October 2024)
[…]
1 Introduction
Molecular dynamics (MD) simulations are a powerful tool for studying the structure, dynamics, and interactions of molecular systems. However, generating conformational ensembles and sampling rare events, e.g., protein-ligand binding, remains challenging due to high computational costs and the complexity of the associated energy landscapes[38]. Interactive molecular dynamics in virtual reality (iMD-VR) has recently emerged as a promising approach to address these challenges by leveraging human intuition during real-time MD simulations within an immersive 3D environment[47]. In iMD-VR, users can directly manipulate and steer molecular systems using natural hand motions, applying forces to drive physically-relevant rare events such as conformational changes and ligand binding/unbinding [27]. This human-in-the-loop approach leverages the human’s innate ability for 3D spatial reasoning and manipulation, enabling a user to intuitively explore complex molecular landscapes. Recent studies have demonstrated the efficacy of iMD-VR in recreating crystallographic binding poses for protein-ligand systems[8] [7] and generating important reactive pathways[44]. These interactive simulations capture valuable conformational data that can be challenging to obtain through conventional MD alone, thus offer new opportunities for applications such as training machine learning and investigating reaction mechanisms.
Imitation learning (IL), or learning from demonstration, is a powerful paradigm in artificial intelligence, enabling machines to acquire new skills by observing and mimicking expert behavior[52]. This approach has been particularly influential in the field of robotics, where it has been used to teach robots complex tasks without the need for explicit programming. By observing human demonstrations, robots can learn to perform a variety of actions, ranging from simple manipulations to complex, multi-step procedures[39][32][13][50]. Learning from observation differs from other types of machine learning such as reinforcement learning, where an explicit reward function needs to be defined in advance or fine-tuned during training. IL offers the ability to learn a mapping of observations to actions done by an expert in demonstrations, making it particularly well-suited to domains for which specifying a reward function is challenging or where human expertise can be leveraged. The versatility of IL has also sparked interest in its application beyond robotics, such as in molecular dynamics, where it could potentially streamline the process of simulating and understanding complex molecular interactions.
IL often requires a large number of demonstrations to effectively learn a policy, especially for complex tasks. Collecting a sufficiently large and diverse dataset of human demonstrations can be challenging for various reasons [50]. One way to think of it is that human behavior is often multi-modal—there are many valid ways to perform a task. Standard imitation learning approaches may average out these modes and learn a sub-optimal policy. Capturing and replicating diverse human behaviors is an open challenge [15], mainly, because robustly capturing all relevant aspects of human demonstrations can be difficult due to sensor limitations, occlusions, etc [21]. This is especially true when collecting data outside of lab settings ‘in the wild’. Therefore there is a growing need for more large-scale, open datasets of human demonstrations on standardized tasks in order to facilitate reproducible research and benchmark imitation learning algorithms[15][48].
Virtual reality (VR) presents a novel and immersive platform for enhancing the capabilities of IL, particularly within the field of iMD-VR where VR is combined with high-performance computing to provide an interactive environment in which researchers can manipulate molecular structures in real-time. The intuitive and engaging nature of VR could revolutionize the way scientists interact with molecular simulations, making it easier to collect data, hypothesize, and test the dynamics of molecular systems[8][42].
This paper aims to explore IL’s current applications in various domains, including key concepts like behavioral cloning and generative adversarial imitation learning. Particular focus is given to its potential in molecular dynamics using VR. We aim to provide a comprehensive review of the existing literature on IL, identify the benefits and challenges associated with its use, and propose innovative ways in which VR could serve as a platform for data creation and collection in MD. By bridging the gap between IL and VR, we hope to open new avenues for research and application in the field of MD, ultimately contributing to advancements in scientific understanding and technological development.
2 Virtual reality for molecular simulations
2.1 Molecular vizualisation
Virtual reality (VR) is revolutionizing the way researchers interact with and visualize molecular structures. VR provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures in way that is not facilitated with traditional 2D interfaces. This can enhance the researcher’s understanding of complex 3D molecular arrangements and interactions, which is essential for enabling research insight. Furthermore, VR can enhance scientific collaboration by providing shared virtual environments, which may even be accessible over the internet, thus enabling collaboration across physical distances.
There exist several programs for the visualization of molecular simulations in VR, e.g., UnityMol[10], and the commercial software Nanome[1]. Nanome provides a collaborative virtual environment in which users can visualize and manipulate molecular structures in stereoscopic 3D. Researchers can analyze the spatial arrangement of molecules, measure distances between atoms, and dock ligands into protein binding pockets using natural hand gestures [18]. Other examples of software include ProteinVR [2] and Molecular Rift [26]. ProteinVR is a web-based application that works across desktop, mobile, and VR platforms, democratizing access to structural biology in 3D. Molecular Rift provides controller-free manipulation of molecules using intuitive hand gestures.
By immersing users in 3D virtual environments, VR enables intuitive exploration of complex biomolecular systems that traditional 2D screens do not facilitate. The application of VR to molecular simulations unlocks several key benefits. First, it provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures. Second, by coupling interactive molecular dynamics with VR, scientists can manipulate molecular systems and observe the effects in real-time, potentially uncovering new mechanistic insights. Third, VR enables collaborative drug design and molecular modeling in shared virtual environments, enhancing scientific teamwork. Finally, the stereoscopic depth perception and wide field of view in VR leads to an enhanced spatial understanding of 3D molecular arrangements.[5]
2.2 Interactive molecular simulations
Recent advancement in computational power and improving performance of graphical processing units has facilitated not only the visualization of molecules in VR, but also real-time interactivity. One prominent example is Narupa, an open-source program developed by Glowacki et al.[14] for performing interactive molecular dynamics in virtual reality (iMD-VR). Using VR controllers, Narupa users can apply forces directly to MD simulations in real-time to drive important chemical events such as ligand binding and conformational changes.
Figure 1: Showing an iMD-VR user docking and undocking the drug oseltamivir from the H7N9 neuraminidase protein [27]
An example of this is demonstrated in Figure 1. The top panel illustrates a researcher using iMD-VR to explore the binding and unbinding pathways of the drug oseltamivir (shown in magenta) with the H7N9 neuraminidase protein. The bottom panel shows snapshots of the molecular system at three timepoints in the simulation: (1) oseltamivir bound to the active site of neuraminidase, (2) the molecular system after the researcher has undocked oseltamivir from the binding pocket of neuraminidase, and (3) the protein-ligand complex, where oseltamivir has been re-docked by the researcher after interactively exploring potential binding modes.
2.3 Data structure in NanoVer
In essence, a molecular simulation is a time series consisting of a set of frames of the atomic positions of a molecular system, which can be viewed frame-by-frame as a ‘trajectory’. In iMD-VR, these trajectories are generated and visualised on-the-fly. Each frame contains information about the system, e.g. the temperature and energy, and the atoms contained within it, e.g. their position and element type. NanoVer (previously known as Narupa) uses a key-value system to store and communicate these frames.
NanoVer streams data in two dictionaries: (a) the frames, containing information about the simulation, such as the atomic positions; and (b) the ‘shared state’, consisting of synchronised information about, e.g., avatar positions and user interactions with the simulation. The NanoVer server can record these data, thus enabling post-hoc analysis and playback of sessions. These recordings can be loaded onto the server, which then sends the recorded streams (synchronised using timestamps) to the clients as if they were real-time simulation streams. In this case, NanoVer acts purely as a molecular visualiser, affording the user control over the position/rotation/scale of the simulation and providing typical playback features (play/pause/restart), though naturally the user cannot apply forces to the molecular system. NanoVer recordings can be imported and analysed using a Python script with the MDanalysis module[33][25].
2.4 Types of molecular simulations
NanoVer provides several molecular simulations that any user can load, or users may import their own OpenMM systems. One of the prototypical examples that we use for NanoVer demonstrations is the simulation of a methane molecule and a carbon nanotube, a molecular system relevant to the study of biomolecular channels that act as molecule-selective filters[17]. In this simulation, players can simulate the action of the nanotube as a biomolecular channel by threading the methane through the nanotube (Figure 2). The molecular system comprises 65 atoms: 60 carbons for the nanotube (labelled C1–C60), and 1 carbon and 4 hydrogens for the methane (labelled C61 and H1–H4). Table 1 shows some example data collected for this task. The output trajectory file ’.traj’ is a binary file containing all data from the molecular simulation. This was converted into a .csv file and the relevant data was filtered for simpler access and processing. The resulting dataframe had 4 columns: atom name, time, coordinates, and user forces. Here ‘name’ is the atom’s label, ‘time’ is the frame index, ‘coordinates’ contains the (x,y,z) positions in nanometers, and ‘forces’ contains the (x,y,z) components of the forces (Fx, Fy, Fz) applied by the user on the specified atom.
Figure 2: Threading Methane through the Nanotube)
Table 1: Dataframe example of the first frame from a recording for the nanotube task
atom name time coordinates user forces
C1 0 [9.725553, 14.941643, 14.158468] [0.0, 0.0, 0.0]
C2 0 [10.063371, 15.170232, 12.954147] [0.0, 0.0, 0.0]
C3 0 [11.367319, 15.154369, 12.419062] [0.0, 0.0, 0.0]
C4 0 [11.99453, 16.465868, 12.049124] [0.0, 0.0, 0.0]
H4 0 [7.0092716, 18.310032, 12.723206] [0.0, 0.0, 0.0]
For analysis purposes, the trajectory of the C61 atom from the methane molecule can be plotted as seen in Figure  3. Each subplot represents a distinct attempt to thread the methane molecule through the nanotube.
Figure 3: Atom C61’s trajectory for Nanotube task
A more sophisticated task is the one shown in Figure 4. This task involves tying a knot using a 17 alanine molecule. The user is presented with the molecule in an untied form and asked to form a knot using the movement of both ends of it. Both given examples are demonstration systems involving only small molecules. The software can be used on larger, more complex molecular systems such as protein systems.
Figure 4: Knot tying task in 17 Alanine molecule
3 Imitation Learning in Agents and Multiagent systems
3.1 Recent works in literature
In recent years, imitation learning in agents and multiagent systems has seen significant advancements. One notable contribution is the introduction of Multi-agent Inverse Factorized Q-learning (MIFQ), a novel algorithm that employs mixing networks to aggregate decentralized Q functions for centralized learning and uses hyper-networks to generate weights for mixing networks [51]. This approach has demonstrated superior performance compared to baseline algorithms in various multi-agent environments, including SMACv2, Gold Miner, and Multi Particle Environments. MIFQ enables efficient and stable learning in cooperative multi-agent settings, and its objective function exhibits convexity within the Q function space under certain conditions.
Another significant development in the field focuses on scaling laws for imitation learning in single-agent games. This research investigates the impact of scaling up model and data size on imitation learning performance, particularly in Atari games and NetHack [28]. By using Behavioral Cloning (BC) to imitate expert policies, the study reveals that imitation learning loss and mean return follow clear power law trends with respect to FLOPs. Importantly, loss and mean return are highly correlated, indicating that improvements in loss predictably translate to improved performance. The research demonstrates that scaling up model and data size can provide significant improvements in agent performance, with the scaled-up approach surpassing prior state-of-the-art by 1.5x in all settings for NetHack.
In the realm of multi-agent systems, the Multi-Agent Adversarial Interaction Priors (MAAIP) approach adapts Multi-Agent Generative Adversarial Imitation Learning (MAGAIL) for modeling interactions between agents [20]. This method introduces new objectives for training the system and models self and opponent observations separately. MAAIP has proven effective for learning interactive behaviors between multiple agents and can be applied to scenarios where agents need to adapt to each other’s actions. This approach demonstrates potential for improving imitation learning in competitive or cooperative multi-agent settings.
These recent advancements collectively highlight the importance of scaling, efficient centralized learning in decentralized execution settings, and modeling agent interactions for improved performance in complex environments. As the field of imitation learning continues to evolve, these contributions pave the way for more sophisticated and effective agent behaviors in both single-agent and multi-agent systems.
Imitation learning has found a wide range of applications in robotics, demonstrating its versatility and effectiveness in enabling robots to perform complex tasks. This section delves into three primary areas where imitation learning has been significantly applied: manipulation tasks, locomotion and navigation, and human-robot interaction.
3.1.1 Manipulation tasks
Manipulation tasks involve robots handling, moving, or altering the state of objects in their environment. In recent years, Imitation learning has been instrumental in teaching robots to perform such tasks with precision and adaptability[31]. For instance, VIOLA [53] a novel IL approach that was implemented and deployed into a real-life robot, outperforms state-of-the-art methods by 45.8% in success rate. This is achieved through the use of a pre-trained vision model which is put into a transformer-like architecture. The authors created a policy to detect task-driven relevant regions for action mapping. Another novel hybrid imitation learning (HIL) framework combines behavior cloning (BC) and state cloning (SC) methods to efficiently learn manipulation tasks like pick-and-place and stacking[16]. This approach has been shown to significantly improve training efficiency and policy flexibility, demonstrating a performance improvement and faster training time compared to pure BC methods. Hua et al. [13] emphasize the efficiency of learning from good samples and the potential for combining reinforcement learning mechanisms to improve the speed and accuracy of imitation learning. They specifically addressed the application of imitation learning in robot manipulation by observing expert demonstrations, which can be generalized to other unseen scenarios.
3.1.2 Locomotion and navigation
In robotics, locomotion and navigation are two fundamental aspects that enable robots to move and operate within their environments effectively. These concepts are crucial for the development of autonomous systems that can perform a wide range of tasks, from simple delivery services to complex exploration missions. Locomotion refers to the various methods that robots use to move from one place to another. This movement can be achieved through different mechanisms, depending on the robot’s design and the environment it is intended to operate in. Navigation involves the process by which a robot determines its position in the environment and plans a path to reach a specific destination. It encompasses several key competencies: Self-Localization; The ability of a robot to establish its own position and orientation within a frame of reference. Path Planning; Once the robot knows its location, path planning involves determining the most efficient or safest route to reach the desired destination. Map-Building and Map Interpretation; For effective navigation, robots often need to construct or utilize maps of their environment. This involves sensing the surroundings to identify obstacles, pathways, and other relevant features, and interpreting this information to make navigation decisions.
In the research paper "Learning to Walk by Steering: Perceptive Quadrupedal Locomotion in Dynamic Environments," Seo et al. [41] introduce PRELUDE, a hierarchical learning framework designed to enhance the navigation and locomotion capabilities of quadrupedal robots in dynamic and cluttered environments. The framework divides the problem into two levels: high-level navigation decision-making and low-level gait generation. The high-level controller is trained using imitation learning from human demonstrations collected with a steerable cart, enabling the robot to acquire complex navigation behaviors. The low-level gait controller is trained through reinforcement learning, allowing the discovery of versatile gait patterns through trial and error. The effectiveness of PRELUDE is demonstrated through simulations and hardware experiments, showing significant improvements over state-of-the-art reinforcement learning methods in terms of success rate and travel distance in various environmental conditions[41]. This work exemplifies the application of imitation learning in robotics, particularly in the development of autonomous systems capable of agile and adaptive movement in real-world scenarios.
3.1.3 Human robot interaction
Human interaction tasks in robotics involve robots engaging in various forms of social interaction and cooperation with humans to achieve shared goals. These tasks encompass direct physical interaction, such as assisting with lifting objects or providing physical therapy, as well as collaborative interaction, where robots and humans work together to complete tasks like assembling products on a manufacturing line. Remote interaction, where humans control or collaborate with robots from a distance, also falls under the umbrella of human-robot interaction tasks. Additionally, robots should be able to learn new tasks from human demonstrations and proactively seek human assistance when needed during task execution. The ultimate goal in human-robot interaction tasks is to achieve natural, efficient, and safe interactions as robots work with humans across various domains. Mehta et al. [23] introduce a learning formalism that unifies approaches for physical human-robot interaction by incorporating demonstrations, corrections, and preferences. It represents a comprehensive approach to learning from human interactions, aiming to improve robot adaptability and performance in collaborative tasks. This framework is designed to learn without making assumptions about the tasks the human wants to teach the robot. The key insight of the paper is that physical human-robot interaction can be a rich source of information for teaching robots, and that by leveraging all available forms of interaction—kinesthetic guidance (demonstrations), adjustments to the robot’s motion (corrections), and evaluative feedback (preferences)—a more robust and flexible learning system can be developed. The authors propose a two-step algorithm that first learns a reward model from scratch by comparing the human’s input to nearby alternatives and then applies constrained optimization to map the learned reward into a robot trajectory. This process is iterative and allows for real-time updates based on the human’s feedback, which can be provided in any order and combination. The approach is validated through simulations and a user study, demonstrating that it can more accurately learn manipulation tasks from physical human interaction than existing baselines, especially when faced with new or unexpected objectives.[23] The paper’s insight emphasizes the importance of a unified learning approach that does not rely on predefined task features or reinforcement learning, thus enabling robots to learn new and unexpected tasks in real-time from physical interaction with humans. This has significant implications for the development of robots that can adapt to a wide range of tasks in shared human-robot environments, such as factories, homes, or healthcare settings, where safety and adaptability are paramount.
3.2 Key concepts and techniques
3.2.1 Behavioral cloning
Behavioral cloning (BC) is a straightforward approach that treats imitation learning as a supervised learning problem[30]. Given a dataset of state-action pairs from expert demonstrations, behavioral cloning directly learns a policy (mapping from states to actions) using regression or classification algorithms[37]. The policy is trained to minimize some loss function between the predicted and demonstrated actions on the training data. Based on the demonstration quality, BC is somewhat simple to implement since no extensive knowledge of the environmental dynamics is required. Being treated as supervised learning, a method that is very well studied, makes training BC algorithms computationally efficient.
Consider a dataset $\mathcal{D}=\{(s_{i},a_{i})\}_{i=1}^{M}$ consisting of state-action pairs collected from an expert policy $\pi^{*}$, where $s_{i}$ represents the state and $a_{i}$ the action taken by the expert in that state. The objective of behavioral cloning is to learn a policy $\hat{\pi}$ that approximates the expert policy $\pi^{*}$ as closely as possible.
The process involves, as seen in Figure 5:
  1. 1.
    Data Collection: Collect a dataset $\mathcal{D}$ of state-action pairs $(s,a)$ by observing an expert performing the task.
  2. 2.
    Learning: Train a model $\hat{\pi}$ on $\mathcal{D}$ to learn the mapping from states to actions. This typically involves minimizing a loss function over the dataset. For discrete action spaces, a common choice is the negative log-likelihood (NLL) loss:
    $\mathcal{L}(\pi,s,a^{*})=-\ln\pi(a^{*}|s)$ (1)
    For continuous action spaces, the mean squared error (MSE) loss is often used:
    $\mathcal{L}(\pi,s,a^{*})=\|\pi(s)-a^{*}\|^{2}$ (2)
  3. 3.
    Policy Output: After training, the learned policy $\hat{\pi}$ can be used to perform the task, ideally replicating the expert’s performance.
Figure 5: Process of Behavioral cloning
3.2.2 Inverse reinforcement learning
Inverse Reinforcement Learning (IRL) is a method used to infer the reward function of an agent by observing its behavior within an environment. It assumes that we observe an agent following an unknown policy $\pi^{*}$ and we want to infer the reward function $R$ that this policy is optimizing. The problem is challenging because there are potentially many reward functions that could explain the observed behavior, making IRL an ill-posed problem. IRL is typically modeled as a Markov Decision Process (MDP) where the goal is to determine what objectives or values the agent is optimizing for, given its observed actions.
To understand IRL, we first need to understand the framework in which it operates, which is the MDP. An MDP is defined by a tuple $(S,A,T,\gamma,R)$:
- $S$ is a set of states. - $A$ is a set of actions. - $T$ is the transition probability matrix, where $T(s^{\prime}|s,a)$ gives the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. - $\gamma$ is the discount factor, which determines the present value of future rewards. - $R$ is the reward function, which assigns a scalar reward to each state (or state-action pair).
A policy $\pi$ is a mapping from states to actions, and the goal in reinforcement learning is to find an optimal policy $\pi^{*}$ that maximizes the expected sum of discounted rewards.
The general approach to IRL involves the following steps, see Figure 6:
  1. 1.
    Collecting Data: Observe the behavior of the expert agent and collect state-action trajectories.
  2. 2.
    Estimating the MDP: Use the collected data to estimate the transition probabilities $T$ and the initial state distribution.
  3. 3.
    Learning the Reward Function: Infer a reward function $R$ that would make the observed behavior appear optimal.
The mathematical formulation of IRL can be described as follows:
  • Given:
    • A set of observed trajectories $\tau=\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$ from an expert policy $\pi^{*}$.
    • An estimated MDP $(S,A,T,\gamma)$ without the reward function.
  • Find:
    • A reward function $R:S\times A\rightarrow\mathbb{R}$ such that the expert policy $\pi^{*}$ is optimal for this reward function.
Figure 6: Process of Inverse Reinforcement Learning (IRL)
One common approach to solving IRL is to use a linear approximation of the reward function, where $R(s,a)=\theta^{T}\phi(s,a)$, and $\phi(s,a)$ is a feature representation of the state-action pair. The parameters $\theta$ are then learned by optimizing a likelihood function or by matching the feature expectations of the expert’s policy. Several algorithms have been proposed for IRL, including:
  • Maximum Entropy IRL: This method assumes that the expert behaves in a way that maximizes entropy, meaning that among all policies that could explain the expert’s behavior, the one that is least committed to unnecessary constraints is chosen.
  • Bayesian IRL: This approach treats the reward function as a random variable and uses Bayesian methods to infer a posterior distribution over the reward function given the observed behavior.
3.2.3 Adversarial imitation learning
IRL algorithms seen previously have a high computational complexity [24][9] since they require the execution of RL in inner loops [11]. Adversarial imitation learning has been proposed as a solution to this computational challenge[12]. Generative Adversarial Imitation Learning (GAIL) is an adversarial imitation learning algorithm that uses the framework of generative adversarial networks (GANs) to directly learn a policy from expert demonstrations, without needing to first learn a reward function as in inverse reinforcement learning[12].
The key idea is to train a generator policy to produce trajectories that are indistinguishable from the expert trajectories, as judged by a discriminator network. This is formulated as a minimax game between the generator and discriminator[12]:
$\min_{\pi}\max_{D}\mathbb{E}_{\pi}[\log D(s,a)]+\mathbb{E}_{\pi_{E}}[\log(1-D(s,a))]-\lambda H(\pi)$ (3)
where:
  • $\pi$ is the generator policy
  • $D$ is the discriminator
  • $\pi_{E}$ is the expert policy
  • $H(\pi)$ is an entropy regularization term
Let $\rho_{\pi}(s,a)$ denote the occupancy measure, i.e. the distribution of states and actions encountered when navigating the environment with policy $\pi$[12].
GAIL seeks a policy whose occupancy measure matches the expert’s:
$\rho_{\pi}(s,a)\approx\rho_{\pi_{E}}(s,a)$ (4)
It can be shown that finding a policy to minimize the Jensen-Shannon divergence between occupancy measures is equivalent to the following:
$\arg\min_{\pi}-H(\pi)+\psi_{GA}(\rho_{\pi}-\rho_{\pi_{E}})$ (5)
where $\psi_{GA}$ is a convex regularizer with the form:
$\psi_{GA}(\rho_{\pi})=\max_{D}\mathbb{E}_{\pi}[\log D(s,a)]+\mathbb{E}_{\pi_{E}}[\log(1-D(s,a))]$ (6)
This leads to the GAIL objective in the first equation. The discriminator $D$ is trained to distinguish expert vs policy state-action pairs, while the policy $\pi$ is trained to maximize the discriminator confusion.
The GAIL algorithm alternates between training the discriminator and taking policy gradient steps:
  1. 1.
    Sample trajectories $\tau_{i}\sim\pi_{\theta_{i}}$ from current policy
  2. 2.
    Update discriminator parameters to maximize: $\mathbb{E}_{\tau_{i}}[\nabla_{w}\log D_{w}(s,a)]+\mathbb{E}_{\tau_{E}}[\nabla_{w}\log(1-D_{w}(s,a))]$
  3. 3.
    Take a policy gradient step using cost function $\log D_{w}(s,a)$, e.g. using TRPO[40] (Trust Region Policy Optimization)
The policy optimization uses a model-free RL algorithm like TRPO, using the discriminator output as the reward signal. Training continues until the policy performs well and the discriminator is unable to distinguish policy and expert state-action pairs.
GAIL leverages the expressive power of GANs to directly imitate expert demonstrations, without needing to recover a reward function explicitly as in inverse RL. The discriminator learns to distinguish expert data, providing a reward signal to optimize the policy to match the expert’s occupancy measure. This enables the imitation of complex behaviors from a relatively small number of demonstrations.
4 Strategies for uses in iMD-VR
Imitation learning in a fully simulated VR environment for interactive molecular dynamics offers several practical advantages over the current approach used in robotics. In molecular dynamics simulations, both training and inference can be conducted entirely within the virtual environment, eliminating the need to bridge the gap between simulation and physical reality. This approach is more practical for several reasons.
Firstly, there is consistency between training and deployment. Unlike robotics, where training occurs in VR but deployment happens in the real world, molecular simulations maintain a consistent virtual environment throughout. This eliminates the "reality gap" that often plagues robotic applications, where policies learned in simulation may not transfer perfectly to real-world scenarios.
Secondly, VR-based molecular dynamics simulations offer superior scalability and data generation capabilities. Researchers can generate vast amounts of training data quickly and efficiently, creating diverse scenarios and interactions without the physical constraints or safety concerns associated with real-world robotic systems. This is particularly valuable for exploring complex molecular systems and interactions that would be difficult or impossible to replicate in physical experiments. Furthermore, the simulated environment allows for precise control over all variables, enabling researchers to isolate specific factors and study their effects on molecular interactions. This level of control is often impossible or impractical in physical robotic setups. Researchers can manipulate individual atoms, adjust environmental conditions, and explore extreme scenarios that would be challenging or dangerous to replicate in the real world.
Cost-effectiveness is another significant advantage. Conducting both training and inference in a virtual environment significantly reduces hardware costs and eliminates the need for expensive robotic equipment. This makes the research more accessible to a broader range of institutions and researchers. Additionally, virtual simulations can be run on cloud-based systems, further reducing infrastructure costs and enabling collaborative research across different locations. Safety and repeatability are also key benefits. Virtual molecular dynamics simulations can explore potentially hazardous or extreme conditions without risking damage to physical equipment or compromising safety. Experiments can be repeated indefinitely with exact precision, facilitating rigorous scientific investigation and enabling researchers to explore a wider range of scenarios than would be possible in physical experiments.
Imitation learning approaches like [53], [16], and learning from good samples could be applied to the task of threading a methane molecule through a carbon nanotube or the knot tying one which we explained previously. For instance, a pre-trained vision model in a transformer-like architecture, similar to VIOLA [53], could be used to detect task-driven relevant regions of the nanotube and methane molecule for precise action mapping. A hybrid approach combining behavior cloning and state cloning, inspired by HIL[16], could efficiently learn the manipulation task of threading the molecule through the nanotube or learn to tie a knot by observing expert demonstrations in VR. In similar implementations of imitation learning, researchers have identified several significant challenges, including covariate shift [29], causal misidentification [6], and the copycat problem [49]. Each of these challenges undermines the effectiveness of imitation learning algorithms, but the literature over the years has proposed innovative solutions to mitigate these issues, leading to more robust and generalizable models.
Covariate shift [29], a prevalent issue where the training data distribution does not match the test data distribution, has been a focal point of concern. This mismatch leads to models that perform well on training data but fail to generalize to new, unseen environments. To combat this, interactive imitation learning (IL) techniques [45] such as DAgger [34] (Dataset Aggregation) have been developed. These methods iteratively refine the training dataset by incorporating data collected under the policy currently being learned, thus aligning the training and test distributions more closely. Furthermore, inverse reinforcement learning (IRL) approaches that focus on learning the underlying reward function from expert demonstrations offer another avenue to address covariate shift. By concentrating on the reward structure rather than directly mimicking actions, these methods aim to achieve better generalization. Additionally, constrained IL [3] introduces constraints into the learning problem to prevent significant deviations from the expert policy, thereby reducing the impact of covariate shift.
Causal misidentification [6], where models learn incorrect causations between observations and actions, poses another challenge. This issue can lead to models that make decisions based on spurious correlations, resulting in suboptimal or incorrect behavior. Researchers have tackled this problem by applying causal inference techniques [46] to distinguish between causally relevant and irrelevant features. Moreover, structured imitation learning methods [35] that incorporate knowledge about the task or environment into the learning process help the model focus on the correct causal relationships, enhancing its decision-making capabilities.
The copycat problem [49], characterized by models mimicking expert actions without understanding the underlying task structure, has also received attention. Solutions such as residual action prediction [4], where the model predicts deviations from the expert’s actions, encourage a deeper understanding of the task dynamics. Temporal regularization, which penalizes large changes in actions over time, further discourages models from blindly copying expert actions, promoting a more nuanced approach to learning from demonstrations.
Madirolas et al. [22] demonstrates that the Wisdom of the Crowd (WOC) effect can be successfully applied to a visuomotor control task of tracing shapes on a touchscreen. The authors show that aggregating the trajectories from a large group of individuals (including children) produces a collective trajectory that is much more accurate than most individual ones. Specifically, the average error of the WOC trajectory was 2-5 times lower than individual trajectories. Importantly, this dramatic improvement required aggregating trajectories from different individuals, not just repeated trials from the same person. The WOC trajectory also outperformed over 99% of the individual trajectories in accuracy. This has important implications for citizen science projects that involve iMD-VR or data collection from large groups of non-expert volunteers. It further demonstrates that for certain types of tasks, aggregating the inputs from many citizen scientists, even if they are not individually highly skilled, can yield results superior to those produced by a single expert.
4.1 Potential applications of IL for iMD-VR
The use of imitation learning for iMD-VR has a range of potential applications. In this section, we identify two such domains: drug discovery and protein engineering, and material design.
4.1.1 Ligand/drug binding to protein
One of the most promising applications of IL for iMD-VR is in computer-aided drug design (CADD)[47]. CADD methods are being used within drug development to reduce the financial and temporal costs associated with the discovery, development and analysis of drug candidates. Sabe et al. [36] reported in 2021 that some form of CADD technique had been used in the development pipeline of more than 70 commercialised drugs. Where the target protein structure is known, MD simulations can be used to shortlist candidate molecules for potential bioactivity by calculating protein-ligand complex stability[43]. However, simulating rare events such as ligand binding using MD still remains a challenge.
iMD-VR has been demonstrated as a human-in-the-loop strategy that leverages the human ability to perform spatial tasks to address the problem of the simulation of ligand binding. Ligand binding is akin to 4-dimensional Tetris, where one 3D shape must fit into another. The difficulty is that these shapes are dynamic and flexible, and that they interact with one another in complex ways. Although this is a difficult task to boil down into an algorithm, humans are able to do this naturally using their spatial intuition combined with their motor skills to perform these types of tasks with minimal training. This is exemplified in a study by Deeks et al.[7], who demonstrated the use of iMD-VR for docking ligands to the main protease of the SARS-CoV-2 virus (the virus responsible for the COVID-19 pandemic). The authors found that iMD-VR experts were able to form docked structures that were in agreement with the crystal structures found experimentally. Another notable study by Deeks et al.[8] found that non-experts could also generate accurate structures of protein-ligand complexes. The authors reported that novice iMD-VR users, many of whom were also not experts in ligand binding, could reliably reproduce experimentally-derived docking poses of flexible ligands with only a short amount of training (<40 minutes in VR). This suggests that IL models could be trained effectively using data gathered from both expert and non-expert users, increasing the size of training sets that leverage human intuition to further sample these non-trivial rare events. IL could greatly enhance the use of iMD-VR in the context of ligand-protein binding by learning from the physically relevant trajectories produced by both experts and non-experts to effectively sample the space of possible binding pathways, leading to a better quantitative understanding of the relative energetics of the docking process that naturally influences the effectiveness of a given drug candidate. This could be extended further to protein engineering by training IL models on iMD-VR-generated datasets of users exploring binding pathways for novel proteins.
4.1.2 Material properties investigation
[…]
Crossley-Lewis et al. [5] also used iMD-VR to examine the transport of the catalytic promoter methyl $n$-hexanoate through the H-ZSM-5 zeolite. In this system, the researchers used iMD-VR to sample the dynamics of the promoter-zeolite system after the rare event of desorption of methyl $n$-hexanoate from a Brønsted acid site within the zeolite framework, an event that is unlikely to be observed on the timescale of a typical unbiased MD simulation.[5] By applying biasing forces to desorb the promoter molecule and pull it into varied positions in the zeolite framework, the researchers were able to investigate the transport dynamics of methyl $n$-hexanoate on accessible timescales, identifying features of zeolite structure relevant to the dynamics of the molecule.[5] To develop this study further, a more quantitative understanding of both the energetics of desorption and the rate of diffusion of the promoter after desorption would be desirable. Once again, this is where IL could assist iMD-VR in a research context: not only could IL help to determine the optimal pathways for desorption, but it could be used to enhance sampling of the subsequent dynamics after such events. These dynamics are necessary to better approximate quantities of interest such as the diffusion coefficient, a measure of the average promoter diffusion (that greatly influences the catalytic efficiency of the material in question[5]), helping guide the search for effective catalysts and promoters.
5 Conclusion
By harnessing the spatial reasoning abilities and domain expertise of researchers performing molecular manipulation tasks in immersive VR environments, rich datasets can be generated to train AI agents via imitation learning techniques. This human-in-the-loop approach shows great promise for efficiently exploring vast conformational spaces of molecular systems. Imitation learning methods have been successfully employed in robotics for learning complex manipulation tasks from demonstrations, and can potentially be adapted to learn policies for interacting with and manipulating molecular structures in iMD-VR. The unique 3D interaction data captured from researchers in iMD-VR systems presents an opportunity to encode human intuition and domain knowledge into the decision-making of AI agents. Key challenges to be addressed include defining appropriate reward functions, and enabling agent generalization across diverse molecular systems. Hybrid approaches combining imitation learning with other AI techniques like reinforcement learning may prove advantageous in this pursuit. While there is still significant work to be done, the intersection of imitation learning and interactive molecular dynamics in VR presents a promising frontier for accelerating scientific discovery and innovation.
[…]
This project was supported by the Xunta de Galicia (Centro de investigación de Galicia accreditation 2019–2022, ED431G-2019/04) and the European Union (European Regional Development Fund—ERDF), as well as by the European Research Council under the European Union’s Horizon 2020 research and innovation program through consolidator grant NANOVR 866559.